# Stability Regularized Cross-Validation

### v1

@@ 4 Conclusion
In this work, we proposed a new approach to hyperparameter selection, namely selecting hyperparameters that minimize a weighted sum of the cross-validation error and the empirical hypothesis stability, with the weight in the weighted sum selected via a nested cross-validation procedure. Across a suite of real-world datasets, our approach improves the out-of-sample MSE by $4\%$ on average for sparse ridge regression and CART, although it does not improve the performance of XGBoost. It also dramatically reduces the amount of out-of-sample disappointment experienced by the user.
Future work could involve developing a tighter bound on the test set error than the bound derived by [11], by controlling more moments of hypothesis stability. Indeed, [5] reports that controlling more moments of a pseudodistribution provides strictly tighter probabilistic guarantees on a quantity. It would also be interesting to investigate the performance of our nested cross-validation procedure across a broader range of machine learning and operations contexts (e.g., neural networks), and to more extensively quantify the role that algorithmic stability plays in out-of-sample performance.
@@ Appendix A Data Generation Process for Heatmaps
Our data generation process for the motivating example follows the data generation process used by [8] and is as follows:
@@
    The rows of the model matrix are generated iid from a $p$-dimensional multivariate Gaussian distribution $\mathcal{N}(\bm{0},\bm{\Sigma})$, where $\Sigma_{ij}=\rho^{|i-j|}$ for all $i,j\in[p]$.
@@
    A “ground-truth" vector $\bm{\beta}_{\text{true}}$ is sampled with exactly $\tau_{\text{true}}$ non-zero coefficients. The position of the non-zero entries is randomly chosen from a uniform distribution, and the value of the non-zero entries is either $1$ or $-1$ with equal probability.
@@
    The response vector is generated as $\bm{y}=\bm{X\beta_{\text{true}}}+\bm{\varepsilon}$, where each $\varepsilon_{i}$ is generated iid from a scaled normal distribution such that $\sqrt{\nu}=\|\bm{X\beta_{\text{true}}}\|_{2}/\|\bm{\varepsilon}\|_{2}$.
@@ Appendix B Heatmaps From Globally Minimizing Five-Fold Cross-Validation Error
Figure 2: Five-fold (left) and test (right) error for varying $\tau$ and $\gamma$, for the overdetermined setting (top, $n=50,p=10$) and an underdetermined setting (bottom, $n=10,p=50$) considered in Figure 1. In the overdetermined setting, the five-fold error is a good estimate of the test error for most values of parameters $(\gamma,\tau)$. In contrast, in the underdetermined setting, the five-fold error is a poor approximation of the test error, and the estimator that minimizes the five-fold error ($\gamma=6.15$, $\tau=5$) significantly disappoint out-of-sample.
B.1 Proof of Theorem 1
@@
The result follows analogously to [11, Theorem 11]; the main novelty in this proof compared to [11, Theorem 11] is the use of a more general notion of hypothesis stability. In particular, Bousquet and Elisseeff [11]’s definition is sufficient to derive the result for leave-one-out, but not for $k$-fold.
Let $\mathcal{R}:=\frac{1}{|\mathcal{S}|}\ell(y_{i},\bm{\beta}(\bm{\theta},\bm{x}_{i}))$ denote the average test set error on an unseen observation, and $\mathcal{R}_{CV}:=\frac{1}{n}\sum_{j\in[k]}h_{j}(\bm{\theta})$ denote the average $k$-fold cross-validation error. Further, let $\mathbb{E}[\ell(A_{\mathcal{S}},z)]$ denote the expected generalization error of a regressor trained on a training set $\mathcal{S}$ and evaluated on an example $z=(\bm{x}_{i},y_{i})$ drawn from the same distribution but not included in the test set. Let $z^{\prime}$ denote an independent draw to $z$, and $\mathcal{S}^{(\mathcal{N}_{j})}$ denote a training set with the $j$th fold of the data omitted.
Then, analogously to [11, Lemma 9], letting $i\neq j$, one can show that
$\mathbb{E}[(\mathcal{R}-\mathcal{R}_{CV})^{2}]\leq$ $\ \mathbb{E}_{\mathcal{S},z,z^{\prime}}[\ell(A_{\mathcal{S}},z)\ell(A_{\mathcal{S}},z^{\prime})]-2\mathbb{E}_{\mathcal{S},z}[\ell(A_{\mathcal{S}},z)\ell(A_{\mathcal{S}^{(\mathcal{N}_{i})}},z_{i})]$ (9)
$\ +\frac{n-n/k}{n}\mathbb{E}_{\mathcal{S}}[\ell(A_{\mathcal{S}^{(\mathcal{N}_{i})}},z_{i})\ell(A_{\mathcal{S}^{(\mathcal{N}_{j})}},z_{j})]+\frac{M}{k}\mathbb{E}_{\mathcal{S}}[\ell(A_{\mathcal{S}^{(\mathcal{N}_{i})}},z_{i})]$
$=\frac{1}{k}\mathbb{E}_{\mathcal{S}}[\ell(A_{\mathcal{S}^{(\mathcal{N}_{i})}},z_{i})(M-\ell(A_{\mathcal{S}^{(\mathcal{N}_{j})}},z_{j})]$ (10)
$+\mathbb{E}_{\mathcal{S},z,z^{\prime}}[\ell(A_{\mathcal{S}},z)\ell(A_{\mathcal{S}},z^{\prime})-\ell(A_{\mathcal{S}},z)\ell(A_{\mathcal{S}^{(\mathcal{N}_{i})}},z_{i})]$
$+\mathbb{E}_{\mathcal{S},z,z^{\prime}}[\ell(A_{\mathcal{S}^{(\mathcal{N}_{i})}},z_{i})\ell(A_{\mathcal{S}^{(\mathcal{N}_{j})}},z_{j})-\ell(A_{\mathcal{S}},z)\ell(A_{\mathcal{S}^{(\mathcal{N}_{i})}},z_{i})]$
$=I_{1}+I_{2}+I_{3},$
where we let $I_{1},I_{2},I_{3}$ stand for the terms in the first, second, and third lines of the right-hand side.
Further, it follows directly from Schwarz’s inequality [see also 11, pp. 522] that $I_{1}\leq\frac{M^{2}}{2k}$ and it follows analogously to [11, pp. 522] that $I_{2}+I_{3}\leq 3M\mu_{h}$. Therefore, we have that
$\mathbb{E}[(\mathcal{R}-\mathcal{R}_{cv})^{2}]\leq\frac{M^{2}}{2k}+3M\mu_{h}.$
Finally, the result follows from Chebyshev’s inequality. ∎
@@ Appendix C Pseudocode for Nested Stability-Regularized Cross-Validation
Input: dataset $\{(\mathbf{x}_{i},y_{i})\}_{i\in[n]}\in\mathcal{X}\times\mathcal{Y}$; number of folds $k$; learning algorithm $\bm{\beta}$; loss function $\ell:\mathcal{Y}\times\mathcal{Y}\rightarrow\mathbb{R}_{+}$; hyper-parameter grid $\mathcal{H}$; stability grid $\Lambda$
Output: outer scores $\{s_{1},\dots,s_{K}\}$, chosen weight $\lambda$, chosen hyper-parameters $h^{\star}$
Randomly partition dataset into $k$ disjoint partitions indexed by $\mathcal{N}_{1},\ldots,\mathcal{N}_{K}$
for $\lambda\in\Lambda$ do
        // Stability weight loop
      for $t\leftarrow 1$ to $k$ do
              // outer loop to evaluate $\lambda$
             $\mathcal{D}_{\text{outer}}\leftarrow\{(\bm{x}_{i},y_{i})\}_{i\in\mathcal{N}_{t}}$
            $\mathcal{D}_{\text{rest}}\leftarrow\{(\bm{x}_{i},y_{i})\}_{i\in[n]\backslash\mathcal{N}_{t}}$
            // inner loop: hyper-parameter search
             foreach $h\in\mathcal{H}$ do
                   $\bm{\beta}^{(\mathcal{N}_{t})}(h)\leftarrow\text{fit\_model}(\mathcal{D}_{\text{rest}},h)$
                   for $t_{2}\leftarrow 1$ to $k:t\neq t_{2}$  do
                         $\mathcal{D}_{\text{train}}\{(\bm{x}_{i},y_{i})\}_{i\in[n]\backslash\mathcal{N}_{t}\cup\mathcal{N}_{t_{2}}}$
                         $\mathcal{D}_{\text{val}}\leftarrow\{(\bm{x}_{i},y_{i})\}_{i\in\mathcal{N}_{t_{2}}}$
                         $\bm{\beta}^{(\mathcal{N}_{t}\cup\mathcal{N}_{t_{2}})}(h)\leftarrow\text{fit\_model}(\mathcal{D}_{\text{train}},h)$
                         $s_{t_{2}}(h)\leftarrow\frac{1}{|\mathcal{N}_{t_{2}}|}\sum_{i\in\mathcal{N}_{t_{2}}}\ell(y_{i},\bm{\beta}^{(\mathcal{N}_{t}\cup\mathcal{N}_{t_{2}})}(h,\bm{x}_{i}))$
                         $\mu(h)\leftarrow\max(\mu(h),\frac{1}{|n-|\mathcal{N}_{t}|}\sum_{i\in[n]\backslash\mathcal{N}_{t}}\left|\ell\left(y_{i},\bm{\beta}^{(\mathcal{N}_{t}\cup\mathcal{N}_{t_{2}})}(h,\bm{x}_{i})\right)-\left(y_{i},\bm{\beta}^{(\mathcal{N}_{t})}(h,\bm{x}_{i})\right)\right|$
                  $\tilde{s}(h)\leftarrow\dfrac{1}{L}\sum_{t_{2}}s_{t_{2}}(h)$
                    // mean inner kCV score
            $h^{\star}_{\lambda}\leftarrow\arg\min_{h\in\mathcal{H}}\tilde{s}(h)+\lambda\cdot\mu(h)$
              // pick best regularized score
             $m_{\lambda,t}\leftarrow\sum_{i\in\mathcal{N}_{t}}\ell(y_{i},\bm{\beta}^{(\mathcal{N}_{t})}(h^{\star}_{\lambda},\bm{x}_{i}))$
              // Evaluate performance of optimized model on remaining fold
$\bar{m}_{\lambda}\leftarrow\dfrac{1}{n}\sum_{t=1}^{k}m_{\lambda,t}$
  // Estimated performance with regularization $\lambda$
$\lambda^{\star}\leftarrow\operatorname*{arg\,min}_{\lambda\in\Lambda}\bar{m}_{\lambda}$
  // Find best $\lambda$ out-of-sample
foreach $h\in\mathcal{H}$ do
       // Now do regularized $k$-fold CV
       $\bm{\beta}(h)\leftarrow\text{fit\_model}(\{(\bm{x}_{i},y_{i})\}_{i\in[n]},h)$
       foreach $t\in[k]$ do
             $\bm{\beta}^{(\mathcal{N}_{t})}(h)\leftarrow\text{fit\_model}(\{(\bm{x}_{i},y_{i})\}_{i\in[n]\backslash\mathcal{N}_{t}},h)$
             $s_{t}(h)\leftarrow\frac{1}{|\mathcal{N}_{t}|}\sum_{i\in\mathcal{N}_{t}}\ell(y_{i},\bm{\beta}^{(\mathcal{N}_{t})}(h,\bm{x}_{i}))$
             $\mu(h)\leftarrow\max(\mu(h),\frac{1}{n}\sum_{i\in[n]}\left|\ell\left(y_{i},\bm{\beta}^{(\mathcal{N}_{t})}(h,\bm{x}_{i})\right)-\left(y_{i},\bm{\beta}(h,\bm{x}_{i})\right)\right|$
      $h^{\star}\leftarrow\arg\min_{h\in\mathcal{H}}\frac{1}{n}\sum_{t\in[k]}s_{t}(h)+\lambda^{\star}\mu(h)$
return $\min_{\lambda\in\Lambda}(\bar{m}_{\lambda}),\lambda^{\star},h^{\star}$

### v2

@@ 4 Conclusion
In this work, we proposed a new approach to hyperparameter selection, namely selecting hyperparameters that minimize a weighted sum of the cross-validation error and the empirical hypothesis stability, with the weight in the weighted sum selected via a nested cross-validation procedure. Across a suite of real-world datasets, our approach improves the out-of-sample MSE by $4\%$ on average for sparse ridge regression and $2\%$ on average for CART, but does not improve XGBoost’s performance. It also dramatically reduces the user’s out-of-sample disappointment.
Future work could involve developing a tighter bound on the test set error than the one derived by (Bousquet & Elisseeff, 2002). It would also be interesting to investigate the performance of our nested cross-validation procedure across a broader range of contexts (e.g., neural networks), and to more extensively quantify the role that algorithmic stability plays in out-of-sample performance.
Impact Statement
This paper presents work aimed at advancing the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
@@ Appendix A Data Generation Process for Heatmaps
Our data generation process for the motivating example follows the data generation process used by (Bertsimas et al., 2020) and is as follows:
@@
    The rows of the model matrix are generated i.i.d. from a $p$-dimensional multivariate Gaussian distribution $\mathcal{N}(\bm{0},\bm{\Sigma})$, where $\Sigma_{ij}=\rho^{|i-j|}$ for all $i,j\in[p]$.
@@
    A “ground-truth” vector $\bm{\beta}_{\text{true}}$ is sampled with exactly $\tau_{\text{true}}$ non-zero coefficients. The position of the non-zero entries is randomly chosen from a uniform distribution, and the value of the non-zero entries is either $1$ or 1-1 with equal probability.
@@
    The response vector is generated as $\bm{y}=\bm{X\beta_{\text{true}}}+\bm{\varepsilon}$, where each $\varepsilon_{i}$ is generated i.i.d. from a scaled normal distribution such that $\sqrt{\nu}=\|\bm{X\beta_{\text{true}}}\|_{2}/\|\bm{\varepsilon}\|_{2}$.
@@ Appendix B Heatmaps From Globally Minimizing Five-Fold Cross-Validation Error
Figure 2: Five-fold (left) and test (right) error for varying $\tau$ and $\gamma$, for the overdetermined setting (top, $n=50,p=10$) and an underdetermined setting (bottom, $n=10,p=50$) considered in Figure 1. In the overdetermined setting, the five-fold error is a good estimate of the test error for most values of parameters $(\gamma,\tau)$. In contrast, in the underdetermined setting, the five-fold error is a poor approximation of the test error, and the estimator that minimizes the five-fold error ($\gamma=6.15$, $\tau=5$) significantly disappoints out-of-sample.
B.1 Proof of Theorem 2.2
@@
The result follows analogously to (Bousquet & Elisseeff, 2002, Theorem 11); the main novelty in this proof compared to (Bousquet & Elisseeff, 2002, Theorem 11) is the use of a more general notion of hypothesis stability. In particular, Bousquet & Elisseeff (2002)’s definition is sufficient to derive the result for leave-one-out cross-validation, but not for $k$-fold cross-validation. Accordingly, we first define all the notation needed for the result, then prove the result.
First, suppose that we fix $\theta\in\Theta$ and let $S=(z_{1},\ldots,z_{n})$ with $z_{i}=(x_{i},y_{i})\stackrel{{\scriptstyle\mathrm{i.i.d.}}}{{\sim}}\mathcal{D}$ be an observation drawn from the stochastic process denoted by $\mathcal{D}$. Further, let $\{\mathcal{N}_{j}\}_{j=1}^{k}$ be a (possibly random, data-independent) partition of $[n]$ into $k$ disjoint folds of equal size $m:=|\mathcal{N}_{j}|=n/k$. Next, for any dataset $T$, write $\beta_{T}(\theta)$ for the predictor trained on $T$ using hyperparameters $\theta$. In particular, define the full-sample predictor $\beta(\theta):=\beta_{S}(\theta)$ and, for each $j\in[k]$, the predictor trained with fold $j$ removed by
$S^{(-j)}:=(z_{i})_{i\notin\mathcal{N}_{j}},\qquad\beta^{(\mathcal{N}_{j})}(\theta)\;:=\;\beta_{S^{(-j)}}(\theta).$
Further, for a point $z=(x,y)$ and a predictor $f$, write $\ell(z,f):=\ell(y,f(x))$. Next, define the (conditional) population risk of the full-sample predictor and the $k$-fold CV estimate by
$R\;:=\;\mathbb{E}_{z\sim\mathcal{D}}\big[\ell(z,\beta(\theta))\big],\qquad R_{\mathrm{CV}}\;:=\;\frac{1}{n}\sum_{j=1}^{k}\sum_{i\in\mathcal{N}_{j}}\ell\!\left(z_{i},\beta^{(\mathcal{N}_{j})}(\theta)\right).$
Further, let $B_{j}:=(z_{i})_{i\in\mathcal{N}_{j}}\in(\mathcal{X}\times\mathcal{Y})^{m}$ denote fold $j$ viewed as a block, and define the block loss
$L(f,B)\;:=\;\frac{1}{m}\sum_{r=1}^{m}\ell(z_{r},f).$
Then $0\leq L(f,B)\leq M$ by assumption and we can rewrite
$R_{\mathrm{CV}}\;=\;\frac{1}{k}\sum_{j=1}^{k}L\!\left(\beta^{(\mathcal{N}_{j})}(\theta),B_{j}\right),\qquad R\;=\;\mathbb{E}_{B\sim\mathcal{D}^{m}}\big[L(\beta(\theta),B)\big].$
We now control the block stability using $\mu_{h}$. In particular, fix $j\in[k]$ and let $B=(z_{1},\ldots,z_{m})\sim\mathcal{D}^{m}$ be independent of $S$. Then
$\mathbb{E}_{B}\Big[\big|L(\beta^{(\mathcal{N}_{j})}(\theta),B)-L(\beta(\theta),B)\big|\Big]$ $\leq\frac{1}{m}\sum_{r=1}^{m}\mathbb{E}_{z_{r}\sim\mathcal{D}}\Big[\big|\ell(z_{r},\beta^{(\mathcal{N}_{j})}(\theta))-\ell(z_{r},\beta(\theta))\big|\Big]$
$=\mathbb{E}_{z\sim\mathcal{D}}\Big[\big|\ell(z,\beta^{(\mathcal{N}_{j})}(\theta))-\ell(z,\beta(\theta))\big|\Big].$
Taking $\max_{j\in[k]}$ and then $\mathbb{E}_{S}$ shows that the stability parameter of the learning rule when evaluated with the block loss $L$ is at most $\mu_{h}$. Next, the standard second-moment bound for stable algorithms (see, e.g., (Bousquet & Elisseeff, 2002, Lemma 9)) applied to the sample of $k$ blocks $(B_{1},\ldots,B_{k})$ with bounded loss $L$ and stability parameter $\leq\mu_{h}$ yields
$\mathbb{E}\big[(R-R_{\mathrm{CV}})^{2}\big]\ \leq\ \frac{M^{2}}{2k}+3M\mu_{h},$
where the expectation is over the draw of $S$ (and any fold-partition randomness). Moreover, Markov’s inequality gives
$\mathbb{P}\big(R-R_{\mathrm{CV}}\geq\varepsilon\big)\leq\mathbb{P}\big((R-R_{\mathrm{CV}})^{2}\geq\varepsilon^{2}\big)\leq\frac{\mathbb{E}[(R-R_{\mathrm{CV}})^{2}]}{\varepsilon^{2}}.$
Setting $\varepsilon:=\sqrt{\frac{\frac{M^{2}}{2k}+3M\mu_{h}}{\delta}}$ and rearranging terms then yields the claim. ∎
@@ Appendix C Pseudocode for Nested Stability-Regularized Cross-Validation
We present the pseudocode in Algorithm 1. Note that when $\lambda=0$, our selection rule corresponds precisely to standard nested cross-validation, where the nesting step is used to better estimate the out-of-sample performance but the same hyperparameters are selected as in $k$-fold cross-validation. The purpose of the nesting scheme is to select $\lambda$, after which the second half of Algorithm 1 is exactly standard $k$-fold cross-validation.
Implementation note: Algorithm 1 is written as if $\Theta$, the set of all hyperparameter combinations, is enumerated (for the sake of clarity). In this case, the models $\bm{\beta}$ need only be computed once for each $\Theta$ and can be reused across all $\lambda$. However, if $\Theta$ is explored adaptively via coordinate descent (as in our experiments), then the nesting step in Algorithm 1 is not redundant.
@@
1:Input: Dataset $\{(\mathbf{x}_{i},y_{i})\}_{i\in[n]}\in\mathcal{X}\times\mathcal{Y}$; number of folds $k$; learning algorithm $\boldsymbol{\beta}$; loss function $\ell:\mathcal{Y}\times\mathcal{Y}\to\mathbb{R}_{+}$; hyper-parameter grid $\bm{\Theta}$; stability-weight grid $\Lambda$.
2:Output: Outer scores $\{m_{\lambda,1},\dots,m_{\lambda,k}\}$; chosen weight $\lambda^{\star}$; chosen hyper-parameters $\bm{\theta}^{\star}$.
3: Randomly partition the dataset into $k$ disjoint folds indexed by $\mathcal{N}_{1},\ldots,\mathcal{N}_{k}$.
4:for $\lambda\in\Lambda$ do
5:  for $t\leftarrow 1$ to $k$ do
6:   $\mathcal{D}_{\text{rest}}\leftarrow\{(\mathbf{x}_{i},y_{i})\}_{i\in[n]\setminus\mathcal{N}_{t}}$
7:   $\triangleright$ inner loop: hyper-parameter search
8:   for all $\bm{\theta}\in\bm{\Theta}$ do
9:    $\mu(\bm{\theta})\leftarrow 0$
10:    $\boldsymbol{\beta}^{(\mathcal{N}_{t})}(\bm{\theta})\leftarrow\textsc{fit\_model}(\mathcal{D}_{\text{rest}},\bm{\theta})$
11:    for $t_{2}\leftarrow 1$ to $k$ do
12:     if $t_{2}\neq t$ then
13:      $\mathcal{D}_{\text{train}}\leftarrow\{(\mathbf{x}_{i},y_{i})\}_{i\in[n]\setminus(\mathcal{N}_{t}\cup\mathcal{N}_{t_{2}})}$
14:      $\mathcal{D}_{\text{val}}\leftarrow\{(\mathbf{x}_{i},y_{i})\}_{i\in\mathcal{N}_{t_{2}}}$
15:      $\boldsymbol{\beta}^{(\mathcal{N}_{t}\cup\mathcal{N}_{t_{2}})}(\bm{\theta})\leftarrow\textsc{fit\_model}(\mathcal{D}_{\text{train}},\bm{\theta})$
16:      $s_{t_{2}}(\bm{\theta})\leftarrow\dfrac{1}{|\mathcal{N}_{t_{2}}|}\sum_{i\in\mathcal{N}_{t_{2}}}\ell\!\left(y_{i},\boldsymbol{\beta}^{(\mathcal{N}_{t}\cup\mathcal{N}_{t_{2}})}(\bm{\theta},\mathbf{x}_{i})\right)$
17:      $\mu(\bm{\theta})\leftarrow\max\!\Biggl(\mu(\bm{\theta}),\;\dfrac{1}{|[n]\setminus\mathcal{N}_{t}|}\sum_{i\in[n]\setminus\mathcal{N}_{t}}\Bigl|\ell\!\left(y_{i},\boldsymbol{\beta}^{(\mathcal{N}_{t}\cup\mathcal{N}_{t_{2}})}(\bm{\theta},\mathbf{x}_{i})\right)-\ell\!\left(y_{i},\boldsymbol{\beta}^{(\mathcal{N}_{t})}(\bm{\theta},\mathbf{x}_{i})\right)\Bigr|\Biggr)$
18:     end if
19:    end for
20:    $\tilde{s}(\bm{\theta})\leftarrow\dfrac{1}{k-1}\sum_{\begin{subarray}{c}t_{2}=1\\ t_{2}\neq t\end{subarray}}^{k}s_{t_{2}}(\bm{\theta})$
21:    $\triangleright$ mean inner CV score
22:   end for
23:   $\bm{\theta}^{\star}_{\lambda}\leftarrow\arg\min_{\bm{\theta}\in\bm{\Theta}}\Bigl(\tilde{s}(\bm{\theta})+\lambda\,\mu(\bm{\theta})\Bigr)$
24:   $\triangleright$ pick best regularized score
25:   $m_{\lambda,t}\leftarrow\dfrac{1}{|\mathcal{N}_{t}|}\sum_{i\in\mathcal{N}_{t}}\ell\!\left(y_{i},\boldsymbol{\beta}^{(\mathcal{N}_{t})}(\bm{\theta}^{\star}_{\lambda},\mathbf{x}_{i})\right)$
26:   $\triangleright$ evaluate on outer fold
27:  end for
28:  $\bar{m}_{\lambda}\leftarrow\dfrac{1}{k}\sum_{t=1}^{k}m_{\lambda,t}$
29:  $\triangleright$ estimated out-of-sample performance at $\lambda$
30:end for
[…]
33:for all $\bm{\theta}\in\bm{\Theta}$ do
34:  $\mu(\bm{\theta})\leftarrow 0$
35:  $\boldsymbol{\beta}(\bm{\theta})\leftarrow\textsc{fit\_model}(\{(\mathbf{x}_{i},y_{i})\}_{i\in[n]},\bm{\theta})$
36:  for $t\leftarrow 1$ to $k$ do
37:   $\boldsymbol{\beta}^{(\mathcal{N}_{t})}(\bm{\theta})\leftarrow\textsc{fit\_model}(\{(\mathbf{x}_{i},y_{i})\}_{i\in[n]\setminus\mathcal{N}_{t}},\bm{\theta})$
38:   $s_{t}(\bm{\theta})\leftarrow\dfrac{1}{|\mathcal{N}_{t}|}\sum_{i\in\mathcal{N}_{t}}\ell\!\left(y_{i},\boldsymbol{\beta}^{(\mathcal{N}_{t})}(\bm{\theta},\mathbf{x}_{i})\right)$
39:   $\mu(\bm{\theta})\leftarrow\max\!\Biggl(\mu(\bm{\theta}),\;\dfrac{1}{n}\sum_{i\in[n]}\Bigl|\ell\!\left(y_{i},\boldsymbol{\beta}^{(\mathcal{N}_{t})}(\bm{\theta},\mathbf{x}_{i})\right)-\ell\!\left(y_{i},\boldsymbol{\beta}(\bm{\theta},\mathbf{x}_{i})\right)\Bigr|\Biggr)$
40:  end for
41:end for
[…]
43:return $\min_{\lambda\in\Lambda}\bar{m}_{\lambda},\;\lambda^{\star},\;\bm{\theta}^{\star}$